what is the faraday’s equivalent of 9.8g of k2cr2o7 ?

Grade:

2 Answers

Abhishek Singh
9 years ago
moles of k2cr2o7 is= 9.8/294=1/30 moles
eqvalents = 1/30*6=0.2 
Abhishek Singh
9 years ago
here 6 is n- factor of k2cr207 (moles of electrons needed to reduce 1 mol. of k2cr2o7)